OpenStudy (anonymous):

6√2 (3√6-7) simplify

OpenStudy (anonymous):



OpenStudy (raden):

first, use the distributive property : a (b - c) = a*b - a*c apply to calculating of 6√2 (3√6 - 7) = ....